OpenStudy (anonymous):

Find an exact solution for Then find the approximate solution to the nearest tenth.

OpenStudy (anonymous):

\[\frac{ \sqrt{5}-1 }{ x } = \frac{ \sqrt{5} }{ 2 }\]